What volume will it occupy at 147°F if the pressure and number of mol are constant?


A sample of oxygen gas has a volume of 2.61 L at 21°C. What volume will it occupy at 147°F if the pressure and number of mol are constant? Enter your answer in the provided box.
Answer #1

Initial volume (V)= 2.61 L Inital temperature to - 21+ 273 = 2quk Final volume (V2) = ? firal temperature = 147°F = (1970f-32
Answer #2

ANSWER :


As per gas law, P V = R T 

R is gas constant and if P remains constant, then,


V / T = constant 

=> V1 / T1 = V2 / T2

Now, 

V1 = 2.61 L

T1 = 21 + 273 = 294 ºK

V2 = ?

T2 = (147 - 32)*5/9 + 273 = 336.89 ºK


So, 


V2 = V1 * T2 / T1 = 2.61 * 336.89 / 294 = 2.99 L (ANSWER).

answered by: Tulsiram Garg